OpenStudy (anonymous):

solve the following system of equations x+3y=5 (1) x=4-3y (2)

OpenStudy (anonymous):

i came up with no solution

OpenStudy (anonymous):

yes no sol

OpenStudy (angela210793):

x+3y=5 (1) x=4-3y (2) replace x in the 1st eq------>4-3y+3y=5 4=5..so ur right

OpenStudy (anonymous):

x+3y=5 x+3y=4